If 2b=a+b thenax+by+c=0 passing through which point

Grade:6

1 Answers

Lucifer Coder
2 years ago
2c = a + b
a + b – 2c = 0
(-a) + (-b) + (2c) = 0
(-a/2) + (-b/2) + (c) = 0
 
Hence, (x,y) = (-1/2,-1/2)